OpenStudy (anonymous):

what is the atomic weight of an element if 4 grams of it contains 2.98 x 10^22 atoms

OpenStudy (anonymous):

find the number of moles first number of molecules = number of moles x Avogadro's constant number of moles = (2.98x10^22)/(6.023x10^23) = 0.0495 moles to find the molar mass number of moles = mass/molar mass molar mass = 4/0.0495 = 80